OpenStudy (anonymous):

What is the value of the expression 15−5 ? A. 1/3125 B. 1/25 C. 25 D. 3125

OpenStudy (anonymous):

@cc_peach

OpenStudy (anonymous):

@cc_peach

OpenStudy (anonymous):

plz i need help

OpenStudy (anonymous):

:'(

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

is it really \[15^{-5}\]??

OpenStudy (anonymous):

yea\s

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok then \[15^{-5}=\frac{1}{15^5}\]

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

\(15^5\)i is a pretty large number so it is probably A

OpenStudy (anonymous):

lol a isn't a large number... its under 1

OpenStudy (triciaal):

if you really need help you would not repost

OpenStudy (misty1212):

actually none of those answers are correct i am going to guess that is it really \[5^{-5}\]

OpenStudy (misty1212):

since \[5^5=3125\] you have \[5^{-5}=\frac{1}{5^5}=\frac{1}{3125}\]

OpenStudy (anonymous):

bout to say... u can't have a negative exponent, but u worked it out

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

What is the value of the expression\[\frac{ 4 }{ 2-2 }\] ? A. ½ B. 1 C. 8 D. 16

OpenStudy (anonymous):

@cc_peach

OpenStudy (anonymous):

u can't divide by zero... 4/2-2 2-2=0 so...??

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):